40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 214 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  7225   Sat Aug 18 17:09:01 2012 SashaUpdateSimulationsC1LSP MEDM Screens Added

C1LSP has been added to the site map. I'll work on filling in the structure some more today and tomorrow (as well as putting up PDH and REFL/AS MEDM screens).

NOTE: Does anyone know how to access channels (or if they're even there) for straight Simulink inputs and outputs (i.e. I have some sort of input, do something to it in the simulink model, then get some output)? I've been trying to add ADC MEDM screens to c1lsp, but channels along the lines of C1LSP-ADC0_0_Analog_Input or C1LSP-ADC0_A0 don't seem to exist.

  7224   Sat Aug 18 03:55:12 2012 YoichiSummaryLSCX-arm locking again

Tonight, I worked on the X-arm locking again. I did not have any significant progress, but observed several issues and will give some suggestions for future work here.

What I did tonight was basically re-alignment of the X-arm (because Rana touched the PZT mirrors for the Y-arm alignment, the X-arm alignment was screwed up). Then I measured the open loop gain. Of course it was almost identical to the one posted in this entry. It reminded myself of how small the phase bubble is. This means we have to finely adjust the gain to set the UGF at the right frequency, i.e. 100Hz. So I decided to do the signal normalization using the TRX power. Using the MC path method described here,  the appropriate normalization coefficient was determined to be 1.6, when the XARM gain is set to 0.05. Using burtgooey, I updated the burt snapshot used by the X-arm restore script.

Now I observed the following things:

When the normalization is used, the lock itself is stable, but the lock acquisition takes loner (i.e. fails more often).

I don't know the exact reason, but here is my guess: Usually, the error signal is divided by the square root of the transmitted power to widen the linear range of the PDH error signal. However, what I'm doing here is dividing the error signal with the power itself, not the sqrt. This might distort the error signal in a not-friendly-for-lock way ? I don't know.

I checked the c1lsc FE code. There seems to be the sqrt(TRX) and sqrt(TRY) signals computed in the code. However, these are not used for the normalization. 

Now, there are two requirements. When dragging the mirrors into the resonance, we want to normalize the error signal with sqrt(TRX). When the mirrors reach the resonance, the gain of the loop must be normalized by TRX. How do we smoothly connect those two states ? Someone should spend some time on this. Maybe I will work on this in Japan.  

We really need a time delay in the filter trigger

The automatic filter trigger is awesome. However, the [0^2:5^2] filter, which is an integrator, takes time to switch on and off. Every time the cavity passes by a resonance, this filter gets turned on and off slowly, giving some large transients. This transient combined with the bad coil balance of ETMX sometimes made the optical lever of ETMX crazy. This can be avoided by turning on this filter a few seconds after the power reaches the threshold. As Rana suggested, we should be able to put an arbitrary time delay to the filter trigger.

Someone should balance the coils

The coil balance of ETMX is bad and causing the above mentioned problem. I tweaked the coil balance by injecting a sinusoidal signal (10Hz) into ETMX pos and trying to minimizing the spectral peak in the optical lever signals. Of course, this is a cheesy work. Someone should put more serious effort on this.

A civilized interferometer should have an auto-alignment capability

After my alignment work, the X-arm power got to about 0.7. (This is probably because the MC transmission power has been low for the past 5 hours or so (attachment 1)).

In anyway, after the cavity locked to the TEM00 mode, the alignment has to be automatically improved by dithering. It is anachronism to sit down and click on the MEDM screen until the power gets big enough.

 

 

Attachment 1: MC_Trans.png
MC_Trans.png
  7223   Sat Aug 18 01:40:09 2012 MashaConfigurationPEMOnline Seismic Noise Classification Widget

I added a widget to the C1PEM_OVERVIEW MEDM screen. The screen shows the nine seismometer channels (GUR1, GUR2, and STS1 X, Y, and Z), the current signal class in dark red, and the overall confidence in the classification, as Rana suggested. The confidence indication thresholds range from 0.1 to 0.9, in intervals of 0.1. Basically, if a signal class is completely dark red, and the other two classes show only white, or, better yet, nothing at all, this means that we have a clear classification. If, however, the other regions have some yellow, or even red indicators, this means that we are not very confident in our signal classification.

Classification_Widget.png

This is a screenshot of the widget. The nine seismometer channels are classifying the signal as quiet, which is good both because it's the middle of the night, and because the nine seismometer signals somehow agree (I'd use the word correspond with one another, but that implies a strong level of coherence..). The confidence is high, seeing as there's little indication in the truck and earthquake regions (none whatsoever in the truck, meaning that the signal, given our classification method, could not possibly be a truck, and some in the earthquake region (below 0.1 of the quiet signal classification strength, however), possibly due to low seismic disturbance).

  7222   Fri Aug 17 18:49:55 2012 ManasaUpdate40m UpgradingOptical layout updated

Quote:

Quote:

ACAD files of the 40m optical layout have been updated as per the vent in Aug 2011.

Files are available at the 40m svn docs-->Upgrade12-->Opt_Layout2011.

 

 To ease the pain of hunting files, optical layout ACAD files have been moved to a new directory 40M_Optical Layout in the repository. Relevant files from directories Upgrade12 and upgrade 08 will be moved to "40M_Optical Layout" very soon and eventually these old directories will be removed. 

Changes mentioned by Koji and Steve have been updated to the files (except for the cable connector which have been added but whose part number has to be found to match accurately with the current layout). The file in the directory should now match the current setup after the last vent Aug 2011.

Let me know if you find any mismatch between the current setup and the layout.

Plans about new installations/reconfiguration during the new vent will be carried out in a separate file.

  7221   Fri Aug 17 18:17:16 2012 MashaConfigurationPEMOnline Seismic Noise Classification - Part 2

As promised in previous e-log, this log is all about the current online seismic noise classification system.

While we had the BLRMS system already in place (which I helped make), Den realized that we would need better filters for the BLRMS channels, as we wanted a strong cut-off, but we also wanted a short step-response so that we could quickly classify seismic signals. Likewise, having a step response which oscillates is also undesirable as this could lead to false classifications of post-truck signal as trucks as a filter adjusts and then dips back down. Thus, after experimenting with many different filters, Den chose to use a combination of

chebyl("LowPass", 1, 1, 0.03)*chebyl("LowPass", 1, 1, 0.03)

as our low-pass filter. The step response and bode plot are below.

LP_RMS_Filter

 The next step was to write C code that would implement the feedforward neural network with my newly generated weights.

Next, I had to implement the code in the c1pem model, and normalize the inputs. Below is an overview of the model, and a close up of the C block section.

GUR1X_Model.png

 GUR1X_Model_Closeup.png

The above close-up includes the process of normalization (dividing by the square of the incoming signal), feeding through the neural network, and classifying.

Each seismometer channel set (GUR1X, GUR1Y, GUR1Z, GUR2X, GUR2Y, GUR2Z, STS1X, STS1Y, STS1Z) now has channels (and corresponding DQ channels) of the following form:

SEIS_CLASS : The class of seismic noise 1.0 means Earthquake, 0.5 means Quiet, and 0.0 means Truck. (There are only these 3 digital values).

SEIS_CLASS_EQ, SEIS_CLASS_TRUCK, SEIS_CLASS_QUIET: These channels represent the confidence of the neural network's classification. The class of the current signal will have an output of 1, where the other two channels will have an output between 0 and 1 representing the ratio of the neural network's output in that class neuron to the output in the classification vector neuron. To simply - suppose the neural network classified an earthquake. Ideally, the neural network output neurons would have the value [1, 0, 0], and SEIS_CLASS would equal 1.0 for earthquake. However, the output neurons probably read something along the lines of [0.9, 0.3, 0.5] - SEIS_CLASS is still 1.0, but SEIS_CLASS_EQ would be 1.0, and SEIS_CLASS_TRUCK would be 0.5 / 0.9 and SEIS_CLASS_QUIET would be 0.3 / 0.9. The lower the other two signals are, the better - this means that we are more confident in our classification.

The MEDM screen for this system (in the RMS system) has the following form for all seismometer channels (this one is GUR1X):

GUR1X_MEDM.png

These are the screens I edited earlier in the summer, with modifications. The bottom filter banks represent the norm of the seismometer signal, which we use to normalize the inputs to the neural network.

Here a close-up of the most important part:

GUR1X_MEDM_CLOSE_UP.png

The orange meter on the right points to the current signal type. Here it reads truck - this is ok because it's the middle of the day, and there are a lot of trucks around. The left side represents our confidence in the signal - the signal is classified as a truck, so the "Truck" bar is saturated. The quiet signal bar is very low, which is good since it means that the neural network thinks that it's definitely not quiet. The earthquake bar has some magnitude, since earthquake signals and trucks have some degree of linear non-separability.

How has this been performing? Firstly, all of the seismometer channels have the same classification readout, which is good. Last night, all of the classes were "quiet", with an "earthquake" which occurred when Den jumped around GUR1 to simulate an EQ. This morning it was on "truck" as expected. The filters are still not fine enough to detect individual trucks, but I will continue to monitor the performance over the coming days.

If anyone has ideas on how better to represent this information, please let me know. This was the first thing that came into my head that would work with my MEDM monitor options, and I'm open to suggestions!

  7220   Fri Aug 17 16:58:06 2012 MashaUpdatePEMOnline Seismic Noise Classification - Part 1

Den and I decided to try to classify seismic signals in the frequency domain rather than the time domain. We looked at amplitude spectral density plots of all of the data in our set, and noted that there were noticeable differences in the frequency domain for midnight quiet, trucks, and earthquakes.

For example, here is the time series of quiet, midnight seismic noise as compared to the seismic noise at the peak of an earthquake - the earthquake signal is noticeably higher in the 1 - 3 Hz region. Likewise, for the truck signal, there are noticeable bumps that arise at 10 and 30 Hz during the peak of the truck's motion due to the resonant frequency of the truck bouncing on its wheels.

noises.png

We investigated this potential means of classification further by considering the linear separability of the power of our signals in various frequency bands. Below is a plot of the power of a normalized signal in the 0.1 - 3.0 Hz region vs. the power of the normalized signal in the 3.0 - 30.0 Hz region - calculated by means of fft and separation of the discrete resulting frequencies (in short, an ideal filter).

Seismic_Signal_Linear_Separability.png

There is rather clear linear separability of the normalized signals in this case, as two lines could potentially be drawn to separate trucks from quiet and earthquake in this case (with a few misclassified points due to quiet - since the lab isn't actually empty and quiet in the middle of the night, and man-made seismic disturbances to occur). The reason we have to normalize our signals lies in the fact that the data set had different gains for various seismometers at different times. Normalization not only allows us to use our data set for training effectively, but it also assures that the online classification, if the online signals are also normalized, will allow for variable seismometer gains in the future and still be able to classify signals.

I looked at the linear separability of our training set using various combinations of frequency bands, and deduced that the current separation in the BLRMS preforms best (coincidentally, since the BLRMS separations are just decades), which meant that we could use the current BLRMS system we have for online classification of seismic noise.

Thus, I built a neural network which performed classification with the following parameters:

- One hidden layer of 20 neurons

- Gradient descent backpropagation with learning parameter mu = 0.175

- Sigmoidal activation functions for each neuron (computationally achieved by a parametrized hyperbola rather than an actual hyper-tangent in order to save on computation time). 

- 5 inputs - the normalized fft^2 of the signal (since the root of a signal doesn't add linearly to 1) in the following frequency regions: 0.1 - 0.3, 0.3 - 1.0, 1.0 - 3.0, 3.0 - 10.0 and 10.0 - 30.0 Hz. Since this division was done through the (frequency, fft value) return in Matlab, the signal was essentially filtered ideally into these frequency bands.

- 3 output neurons representing an output vector, with desired output vectors of [1, 0, 0] for earthquake, [0, 1, 0] for truck, and [0, 0, 1] for quiet.

- 1,600,000 training epochs (batch backpropagation on all of the data)

Below is the best learning curve for this network, representing the total amount of inputs misclassified out of 224. The best result achieved was 30 misclassified signals out of 224. Obviously this is not ideal, but our data is not totally linearly separable. This could, however, be reduced with further iterations, but given the close to 0 slope of the learning curve between iteration number 1,000,000 and number 1,500,000, this could take a very long time.

 

3_Output_Learning_Curve.png

Thus, I trained the network, generated the weight vectors and optimal activation function parameters, and was ready to implement a feed-forward neural network (with no online training). My next e-log (Part 2) will be about this system and will be posted shortly.

Attachment 1: Earthquake_Quiet_PSD.png
Earthquake_Quiet_PSD.png
Attachment 2: Truck_Signal_Progression.png
Truck_Signal_Progression.png
Attachment 3: Seismic_Signal_Linear_Separability.png
Seismic_Signal_Linear_Separability.png
Attachment 4: 3_Output_Learning_Curve.png
3_Output_Learning_Curve.png
Attachment 5: Earthquake_Quiet_PSD.png
Earthquake_Quiet_PSD.png
Attachment 6: Earthquake_Quiet_PSD.png
Earthquake_Quiet_PSD.png
Attachment 7: Truck_Signal_Progression.png
Truck_Signal_Progression.png
  7219   Fri Aug 17 14:45:51 2012 ranaUpdateCamerasvideo cameras in the dark

 The problem with the glow on the ETMY face is due to the red light being scattered off of the optical table from the HeNe laser for the OL. Why is the red light hitting the table?

One way to fix the problem for the camera image is to insert a long pass filter (if Steve can find one).

 Edmund Optics: NT62-874

 Edmund Optics: NT65-731

Edmund Optics: NT32-759 

  7218   Fri Aug 17 12:47:30 2012 SashaUpdateSimulationsThe SimPlant Saga CONTINUES

Quote:

THE GOOD: SimPlants ITMX and ETMX are officially ONLINE. Damping has been instituted in both, with varying degrees of success (see Attachment 1). An overview screen for the SimPlants is up (under XARM_Overview in the sitemap - you can ignore the seperate screens for ETMX and ITMX for now, I'll remove them later), C1LSP will be online/functional by Monday.

The super high low-frequency noise in my simPlant is from seismic noise and having a DC response of 1, so that the seismic noise at low frequencies is just passed as is and then amplified along with everything else in the m --> counts conversion. Not quite sure how to deal with this except by NOT having a DC response of 1 (which it technically doesn't have when you do the algebra - Rana said that "it made sense" for the optic to have unity gain at low frequencies, but the behavior is not matching up with reality).

THE BAD: It looks like the ITMX Switch from Reality to simPlant doesn't work (or some of the signals aren't getting switched). When switching from reality to simulation, it looks like the control system is receiving signals from the SimPlant, but is transmitting them to the real thing. As a result, when you flip the switch from reality to sim, ITMX goes seriously crazy and starts slamming back and forth against the stop. REALLY NOT GOOD. As soon as I saw what was going on, I turned back to reality and flipped the watch dogs on (YES THEY WERE OFF). I'll investigate the connections between the plant and control system some more in the morning (i.e. later today) (this is also probably what is causing the "lost connections" in c1sup/sus we can see in the machine status screen).

 Problem with ITMX solved! The ITMX block in c1sup hadn't been tagged as "top_names". I rebuilt and installed the model, and there are no longer lost connections, :D

  7217   Fri Aug 17 10:38:15 2012 SteveUpdateCamerasvideo cameras in the dark
> I used the LED illuminations at ETMX and BS yesterday for a tour.
> I am afraid that I left them on.

It was turned off before the picture was taken.
All LED illuminations were turned off. I checked them a few times.
  7216   Fri Aug 17 09:34:27 2012 KojiUpdateCamerasvideo cameras in the dark
I used the LED illuminations at ETMX and BS yesterday for a tour.
I am afraid that I left them on.
  7215   Fri Aug 17 08:33:46 2012 SteveUpdateCamerasvideo cameras in the dark

Quote:

 I optimized the TM views with illuminator light on quad1  It actually looks better there.

I'll post a dark-  OSEM light only in jpg tomorrow.  ETMY camera is malfunctioning in dark condition now.

 

ALL  illuminator lighting are off. ITMX and ETMY looks back lighted. I will check on their apertures.

In order to focus on 1064 resonant spots I tried to restore and align the arms  by script. I only got flashes.

Attachment 1: dark.jpg
dark.jpg
  7214   Fri Aug 17 05:29:04 2012 YoichiConfigurationComputer Scripts / ProgramsC1configure scripts

 I noticed that the IFO restore scripts have some problems. They use burt request files to store and restore the settings. However, the request files contain old channel names.

Especially channels with _TRIG_THRES_ON/OFF are now _TRIG_THRESH_ON/OFF, note the extra "H".

These scripts reside in /opt/rtcds/caltech/c1/burt/c1ifoconfigure/.

I fixed the PRMI_SBres and MI scripts. Someone should fix all other files.

  7213   Fri Aug 17 04:54:01 2012 Yoichi, KojiSummaryLSCPRMI Locking

 To taste the strangeness of the current 40m PRC, I locked the PRMI with the guide of Koji.

We first aligned MICH by mostly tweaking ITMX, assuming that ITMY is in a good place as the Y-arm locks. MICH lock was stable.

Then we restored the IFO to the PRM_SBres mode. With a bit of alignment work on PRM and gain tweaking, the PRMI locked.

Yes, the beam spots look UGLY !

Also the PRMI was not so stable. Especially, when the alignment fluctuates, the optical gain changes and the loop becomes temporarily unstable. We took POP_DC as the guide for the gain change and normalized the PRCL error signal with it. To do this smoothly, we first changed the input matrix to route the PRCL error signal, which is REFL33_I, so that the signal also goes to the MC filter bank. Then with dtt, we monitored the spectra of the PRCL_IN1 and MC_IN1. We tweaked the value of the element in the normalization matrix for the MC path until the two spectra look the same (at this moment, the normalizing factor for the PRCL path was still zero). During this process, we noticed that the MC path signal (normalized by POP_DC) is noisier at above 500Hz. This was because the POP_DC has a large noise at high frequencies. So we put a low pass filter (100Hz 2nd order Butterworth) to the POP_DC filter bank to reduce the noise. Then the two spectra looked almost the same. The correct normalization factor found in this way was 0.03. So we put this number in the normalization matrix for PRCL. It did not break the PRMI lock.

 

After the normalization is turned on, the PRMI lock became somewhat more stable. However, the POP_DC was still fluctuating a lot, especially when the alignment is good. So I made a boost filter: 5Hz pole Q=2, 15Hz zero Q=1.5. I also made this filter automatically triggered when the PRMI is locked. This made the PRMI lock acquisition quicker. However, still the POP_DC fluctuation is large. It seems that the alignment of PRC is really fluctuating a lot.

 The current UGF of PRMI is about 150Hz with the phase margin over 50deg.

 

 

 

 

Attachment 1: AS_1029238601.jpg
AS_1029238601.jpg
Attachment 2: POP_1029238616.jpg
POP_1029238616.jpg
Attachment 3: REFL_1029238629.jpg
REFL_1029238629.jpg
Attachment 4: PRMI-OPLG.png
PRMI-OPLG.png
  7212   Fri Aug 17 04:13:45 2012 SashaUpdateSimulationsThe SimPlant Saga CONTINUES

THE GOOD: SimPlants ITMX and ETMX are officially ONLINE. Damping has been instituted in both, with varying degrees of success (see Attachment 1). An overview screen for the SimPlants is up (under XARM_Overview in the sitemap - you can ignore the seperate screens for ETMX and ITMX for now, I'll remove them later), C1LSP will be online/functional by Monday.

The super high low-frequency noise in my simPlant is from seismic noise and having a DC response of 1, so that the seismic noise at low frequencies is just passed as is and then amplified along with everything else in the m --> counts conversion. Not quite sure how to deal with this except by NOT having a DC response of 1 (which it technically doesn't have when you do the algebra - Rana said that "it made sense" for the optic to have unity gain at low frequencies, but the behavior is not matching up with reality).

THE BAD: It looks like the ITMX Switch from Reality to simPlant doesn't work (or some of the signals aren't getting switched). When switching from reality to simulation, it looks like the control system is receiving signals from the SimPlant, but is transmitting them to the real thing. As a result, when you flip the switch from reality to sim, ITMX goes seriously crazy and starts slamming back and forth against the stop. REALLY NOT GOOD. As soon as I saw what was going on, I turned back to reality and flipped the watch dogs on (YES THEY WERE OFF). I'll investigate the connections between the plant and control system some more in the morning (i.e. later today) (this is also probably what is causing the "lost connections" in c1sup/sus we can see in the machine status screen).

Attachment 1: plantvreality.jpg
plantvreality.jpg
  7211   Fri Aug 17 00:16:30 2012 EricSummaryLSCYARM Calibration

I modified my Simulink model of the YARM to match the new filter modules Rana installed on YARM. I also scaled the open loop transfer function of the model to fit the measured open loop transfer function at the unity gain frequency, as shown in the figure below. From this I produced the length response function correctly scaled, also shown below.  Then I applied the calibration factor to the YARM data measured in /users/Templates/Y-Arm_120815.xml. Both the uncalibrated and calibrated spectra are included below.

 

 

Attachment 1: olg_model_meas.png
olg_model_meas.png
Attachment 2: length_response_model.png
length_response_model.png
Attachment 3: yarm_uncal_power_spec.pdf
yarm_uncal_power_spec.pdf
Attachment 4: yarm_cal_power_spec.pdf
yarm_cal_power_spec.pdf
  7210   Thu Aug 16 20:18:39 2012 YaakovUpdateSTACISInput for feedforward/feedback in the STACIS

Below is the bottom view of the geophone preamplifier and controller for the STACIS. It slides into the upper part of the STACIS, under the blue platform. The geophone signal goes in the bottom left, gets amplified, filtered, and otherwise pampered, and goes out from the bottom right. From there it goes on to the high voltage amplifier, and finally to the PZT stacks. Below right is a closer view of the output port for the preamplifier, top and bottom.

SAM_0256.JPGSAM_0259.JPGSAM_0258.JPG

I suggest de-soldering and bending up the pins that carry the geophone signal (so the signals don't go directly to the high voltage amplifier), and adding the circuit below between the preamp and amplifier. The preamp connector is still attached to the high voltage amplifier connector in this setup, only the geophone signal pins are disconnected.

x-chip.png

More on the circuit and its placement:

The first op-amp is a summing junction, and the second is just a unity gain inverter so that signal doesn't go into the high voltage amplifier inverted. I tested this with the breadboard, and it seems to work fine (amplitudes of two signals add, no obvious distortion). The switches allow you to choose local feedback, external feedforward, or both.

The geo input will be wires from the preamp (soldered to where the pins used to go), and the external input will be BNC cables, with the source probably a DAC. The output will go to the bent up pins that used to be connected to the preamp (they go into the high voltage amplifier). This circuit can sit outside of the STACIS- there is a place to feed wires in and out right near where the preamplifier sits. For power, it can use the STACIS preamp supply, which is +/- 15V. The resistors I used in the breadboard test were 10 kOhm, and the op-amp I used was LT1012 (whose noise should be less than either input, see eLog 7190).

This is visually represented below, with the preamp pin diagram corresponding to the soldering points with the preamp upside down (top right picture):

SAM_0266.JPGSAM_0265.JPG

 

  7209   Thu Aug 16 20:04:46 2012 YaakovUpdatePEMTransfer functions of seismic stacks, differential motion of test mass

 

 I made the plots a little nicer and added new sensor noises (from Brian Lantz's scripts and measurements). Click to enlarge.

stack_tf.pngstack_tf.figmass_diff_motion.pngmass_diff_motion.fig

 sensor_noises_diff.pngsensor_noises_diff.fig

The last plot shows that these other sensors' noises are lower than the differential ground motion below 1 Hz.  Though 3 seismometers per STACIS is impractical, this shows that such seismometers could be used as feedforward sensors and provide isolation against differential ground motion. At these noise levels, the noise of the high voltage amplifier circuit in the STACIS would probably be the limiting factor.

  7208   Thu Aug 16 19:12:30 2012 JenneUpdateLockingLong arm lock stretches

YARM_awesome_lock_stretch.png

After Rana and Yoichi tweaked the arm locking filters, we have had some pretty awesome lock stretches. 5-day minute trend.

  7207   Thu Aug 16 18:59:18 2012 JenneUpdateSAFETYEmergency Exit door unlocked again

Koji just found the emergency exit door unlocked again.  NOT GOOD.

We have determined that if you use the emergency door to enter the lab, it leaves the door unlocked, unless you go back outside and deliberately lock it.  This means that someone has been using the emergency exit as a regular entrance. 

It's fine to leave by that door, but you should make a habit of entering through the regular door.  Using the back door as an entrance is a special case situation, when they have the main door blocked.

  7206   Thu Aug 16 17:28:51 2012 ManasaConfigurationIOOMC trans optics configured

Quote:

Quote:

Quote:

Quote:

Quote:

  The PDA255 is a good ringdown detector - Steve can find one in the 40m if you ask him nicely.

 We found a PDA255 but it doesn't seem to work. I am not sure if that is one you are mentioning...but I'll ask Steve tomorrow!

 I double checked the PDA255 found at the 40m and it is broken/bad. Also there was no success hunting PDs at Bridge. So the MC trans is still in the same configuration. Nothing has changed. I'll try doing ringdown measurements with PDA400 today.

Can you explain more what "broken/bad" means?  Is there no signal?  Is it noisy?  Glitch?  etc.

 The PD saturates the oscilloscope just by switching on the power; with no real signal at all. But Steve helped locating a PD that is not being used at the AP table. So I will check it and replace the current one if it works!

Koji opened up the PD and found that the screw connecting the PD to the pole was doing an additional job as well; connecting the power cable to the PD output in the inside. The PD is now fixed! Yippie...we have two PDA255 s at 40m now!!

  7205   Thu Aug 16 16:44:55 2012 ManasaConfigurationIOOPD from AP table removed

Quote:

The PD (pda255) at the AP table, close to the MC refl , which Steve mentioned to be not in use, has been removed from the table for testing.

 The PD installed at MC trans to make ringdown measurements has been replaced with the above PDA255. 

  7204   Thu Aug 16 13:49:33 2012 YaakovUpdatePEMTransfer functions of seismic stacks, differential motion of test mass

I estimated the transfer function of the seismic stacks using a rough model I made based on the LIGO document LIGO T000058 -00. I used a Q of 3.3 for the viton springs, and resonant frequencies of 2.3, 7.5, 15, and 22 Hz (measured in that document for the horizontal motion). I multiplied the simple mass-spring transfer function four times for each layer of metal/spring, with the respective resonant frequency for each. The pendulum suspending the test masses has a resonant frequency of 0.74 and a Q of 3, according to the same document.

stack_tf.pngstack_tf.fig

When I multiply the net transfer function (pendulum included, the green line above) by the differential motion of the x arm that I measured in eLog 7186, I find the differential motion of the test mass (NOTE: I converted the differential motion to displacement by multiplying by (1/2*pi*f)).

mass_diff_motion.pngmass_diff_motion.fig

It agrees within an order of magnitude to the seismic wall from the displacement noise spectrum hanging above the control room computers.

Finally, I looked at how the geophone and accelerometer noise spectra looked compared to the ground differential motion (any STACIS sensor signal will also be multiplied by the stack/pendulum transfer function, so I'm comparing to the differential motion before it goes through the chamber). Below about 1 Hz, it is clear from the plot below that the STACIS could never be of any benefit, even with accelerometers rather than geophones as the feedback sensors.

noise_v_diff_motion.pngnoise_v_diff_motion.fig

Attachment 1: stack_tf.png
stack_tf.png
Attachment 5: stack_tf.fig
  7203   Thu Aug 16 13:04:36 2012 LizSummaryComputer Scripts / ProgramsDaily Summary Details

I just wrote a short description of how to run the daily summary pages and the configuration process for making changes to the site.  It can be found in /users/public_html/40m-summary and is named README.txt.  If I need to clarify anything, please let me know!  The configuration process should be relatively straightforward, so it will be easy to add plots or change them when there are changes at the 40 meter.

  7202   Thu Aug 16 05:08:38 2012 YoichiUpdateIOOMC Servo Transfer Function Measurements

Quote:

 Also the MC board includes many generic filter stages to customize the frequency response. I will open the MC board box to examine what is actually implemented on the board. 

 I took out the MC board. Unfortunately, most of the components are surface mounted. So the values of the capacitors are not legible.

I will try my best to guess what is implemented on the board.

Attachment 1: MCBoard1.JPG
MCBoard1.JPG
Attachment 2: MCBoard2.JPG
MCBoard2.JPG
  7201   Thu Aug 16 01:52:52 2012 YoichiUpdateIOOMC Servo Transfer Function Measurements

Quote:

Last night, I took a bunch of TFs with this method. Now I'm analyzing the data to recover the overall gain G. I will post the results later.

 I calculated the MC open loop transfer function with the combination method. For that, I made a circuit model of the MC board (from the input to the output). The transfer function of this circuit is calculated by SPICE (attachment1). Then it is multiplied by the measured transfer function from the output of the MC board to the input of the MC board (attachment 2) to get the overall transfer function.

The result is shown in the attachment 3. The blue curve is the OPLTF measured with the traditional method. The red curve is the combination method described above. There are some discrepancies between the two curves. The ratio of the two curves (Traditional)/(Combination) is plotted in attachment 4. It seems there is a pole(s) missing from my model of the MC board at around 1MHz. This may come from the omitted op-amps in the MC board model (there are so many op-amps which have flat responses below 1MHz and I omitted most of those). Also the MC board includes many generic filter stages to customize the frequency response. I will open the MC board box to examine what is actually implemented on the board. 

At low frequencies, the two curves are similar but the slope is still different.

I also had to add 83dB of gain to the combined TF to match with the traditional one. I will check where does it come from.

The MC board model (Altium project) is attached as attachment 6. The schematic is attachment 5.

Attachment 1: MC_Board_TF.png
MC_Board_TF.png
Attachment 2: OPTG.png
OPTG.png
Attachment 3: OPLG.png
OPLG.png
Attachment 4: Difference.png
Difference.png
Attachment 5: MC_Board.pdf
MC_Board.pdf
Attachment 6: MC_Board.zip
  7200   Wed Aug 15 20:53:48 2012 ManasaUpdateIOORingdown measurements

Quote:

Quote:

While I thought that the bumps observed at the end of the ringdown might be because of the cavity trying to lock itself, Jan commented that they have always existed in these measurements and their source is not known yet.

What I meant to say was that in all ringdown measurements that we observed today, the bumps were consistently part the ringdown, and that I have no explanation for the bumps. It should also be mentioned that fitting the bumpy part of the ringdown instead (we used the clean first 10us), the ringdown time is about twice as high. In either case, the ringdown time is significantly smaller than we have seen in documents about previous measurements.

We (basically I) also made one error when producing the plots. The yaxis label of the semi-logarithmic plot should be log(...), not log10(...).

 I thought about  why we do not find any bumps beyond the exponential fall. Could it be because we neglected fluctuations of voltage in the negative direction and plotted the absolute values?

  7199   Wed Aug 15 20:15:51 2012 JanUpdateIOORingdown measurements

Quote:

While I thought that the bumps observed at the end of the ringdown might be because of the cavity trying to lock itself, Jan commented that they have always existed in these measurements and their source is not known yet.

What I meant to say was that in all ringdown measurements that we observed today, the bumps were consistently part the ringdown, and that I have no explanation for the bumps. It should also be mentioned that fitting the bumpy part of the ringdown instead (we used the clean first 10us), the ringdown time is about twice as high. In either case, the ringdown time is significantly smaller than we have seen in documents about previous measurements.

We (basically I) also made one error when producing the plots. The yaxis label of the semi-logarithmic plot should be log(...), not log10(...).

  7198   Wed Aug 15 18:56:46 2012 YoichiUpdateIOOMC Servo Transfer Function Measurements

 I started working on the characterization of the MC servo.

The current MC servo topology is shown in the figure attached along with a simplified schematic diagram of the MC board. 

A usual way to measure the open loop gain of this servo is to inject a signal from, say, EXCA of the MC board and measure the transfer function from TP2A to TP1A. It works OK at frequencies around the UGF. The second attachment is the OPLTF measured in this way with the Agilent 4395A. The UGF is about 100kHz with the phase margin of 40 to 50 deg. 

Now we have two issues here. First, I expected the UGF to be more than 100kHz, like 300kHz or so. The phase babble is peaked around 100kHz. According to our old measurement (http://nodus.ligo.caltech.edu:8080/40m/1431) the phase babble peak was at a much higher frequency when the FSS was using the reference cavity. One reason could be that the MC is located much farther from the laser than the reference cavity, so that there is some phase lag caused by the time delay. I will make a model of the MC servo system later to check this theory.

The second issue is that, as you can see in the plot, the OPLTF measurement becomes noisy at lower frequencies. With 4395A, which has the minimum IFBW of 2Hz, OPLTF measurement below 10kHz was impossible with the traditional method. We could use SR785 with a long averaging time to improve the SNR, but it requires a patience which I don't have.

The measurement becomes difficult at low frequencies because the loop gain is too high. When the open loop gain (G) is high, the injected signal (x) from EXCA is immediately suppressed by a factor of 1/(1+G) at TP2A. This makes the injected signal hidden in other noises at TP2A.

How do we solve this problem ? Let's consider a simple servo model shown in the third attachment. A traditional OPLTF measurement is done by injecting a signal from EXC port and measuring the TF from TP2 to TP1. The problem was that at TP2, the signal is attenuated by 1/(1+G1*G2), which is too much when G (=G1*G2) is large. However, at TP3, the attenuated signal is amplified by G1. So the injected signal x becomes x*G1/(1+G) at TP3. If G1's contribution to the overall gain G is large enough,  the signal at TP3 is not so small. Then we can easily measure G2 using TP3 and TP1. In a typical situation, G1 is the transfer function of the electric circuits, which we can know either from standalone measurements or from model calculations, and G2 is an interferometer response, which we want to measure. So, by combining the knowledge of G1 and the measurement of G2, we can obtain the overall loop gain G even at lower frequencies.

 The final attachment shows an example of the measurement of G2. In our case, this is the transfer function from MC_Out_Mon to Q-Mon (see the first attachment) . G1 is the transfer function of the MC board. Since G1 is large at low frequencies, we can measure G2 down to 100Hz with a reasonable integration time (10000 cycles per point).

Last night, I took a bunch of TFs with this method. Now I'm analyzing the data to recover the overall gain G. I will post the results later.

Attachment 1: MC-Diagram.png
MC-Diagram.png
Attachment 2: OPLG-10kHz-1MHz.png
OPLG-10kHz-1MHz.png
Attachment 3: SimpleServoDiagram.png
SimpleServoDiagram.png
Attachment 4: OPTG-100Hz-1kHz.png
OPTG-100Hz-1kHz.png
  7197   Wed Aug 15 17:23:22 2012 jamieUpdateCDSfront end IOP models changed to reflect actual physical hardware

As Rolf pointed out when he was here yesterday, all of our IOPs are filled with parts for ADCs and DACs that don't actually exist in the system.  This was causing needless module error messages and IOP GDS screens that were full of red indicators.  All the IOP models were identically stuffed with 9 ADC parts, 8 DAC parts, and 4 BO parts, even though none of the actual front end IO chassis had physical configurations even remotely like that.  This was probably not causing any particular malfunctions, but it's not right nonetheless.

I went through each IOP, c1x0{1-5}, and changed them to reflect the actual physical hardware in those systems.  I have committed these changes to the svn, but I haven't rebuilt the models yet.  I'll need to be able to restart all models to test the changes, so I'm going to wait until we have a quiet time, probably next week.

  7196   Wed Aug 15 17:17:58 2012 Manasa, JanUpdateIOORingdown measurements

Finally ringdown at IMC conquered and oopsie that came out so clean!

The finesse of the cavity from the current ringdown measurement, F= 453, differs from the measurements made in the document dated 10/1/02 on dcc...not sure if things have changed since then.

While I thought that the bumps observed at the end of the ringdown might be because of the cavity trying to lock itself, Jan commented that they have always existed in these measurements and their source is not known yet.

Ringdown_815.jpg

  7195   Wed Aug 15 16:29:59 2012 Eric SummaryGeneralSURF Update

This week I took more data for the calibration of YARM. The summary of measurements taken is:

1. Peak-to-peak on Michelson
2. Michelson open loop
3. Excite ITMY and measure on AS55_Q_ERR
4. Excite ITMY and measure on POY11_I_ERR
5. Excite ETMY and measure on POY11_I_ERR
6. YARM open loop

Then I worked on comparing these measurements to the Simulink model of the interferometer control loop. The measured open loop transfer function of the YARM matched well with the model above about 20 Hz, after the gain was scaled properly to fit the data. Next is to fit the length response function of the model and the measurements, and then use DTT to calibrate the arm cavity's power spectrum.

  7194   Wed Aug 15 16:01:47 2012 steveUpdateGenerallarger optical tables at the ends ?

The drawing of the 4' x 2'  table cover can be seen at entry  #6190 The new proposed wall #7106  The yellow acrylic would be ~ 0.25" thick and it will be the inside. It is not shown on the drawing.

Question remaining: should get a larger table 4' x 3' as outlined by red lines and make new cover to fit this

The oplev beam path needs larger incident angle to get in and out of the chamber: REMOVE BOTTLENECK for easy traffic

Moving the existing table closer to ETMY chamber - as Jamie suggested-  would help but there is no room for this solution.

The larger table solve this issue and leave more room for initial pointing, arm transmitted and future experiments.

Other benefits: no tube to make between table and chamber. It is easier to make the the larger box air tight.

The new isolation box with feed through, cover, seals will cost $4-5K ea

 

 

Attachment 1: bottleneck.jpg
bottleneck.jpg
  7193   Wed Aug 15 13:24:12 2012 DenUpdateCDSRFM -> OAF

Transmission of signals between RFM and OAF is bad again. Now we do not see any errors in IPC_ERR monitors so models think that they get all data but the data is wrong

oaf.png

  7192   Wed Aug 15 13:23:34 2012 LizSummaryComputer Scripts / ProgramsLast Weekly Update

Over the past week I have been continuing to finalize the daily summary pages, attempting to keep the total run time under half an hour so that they can be run frequently.  I have had many hang ups with the spectrograms and am currently using second trends (with this method, the entire script takes 15 minutes to run).  I also have a backup method that takes 3 minutes of data for every 12 minutes, but could not implement any interpolation correctly.  This might be a future focus, or the summary pages could be configured to run in parallel and full data for the spectrograms can be used.  I configured Steve's tab to include one page of images and one page of plots and fixed the scripts so that it corrects for daylight savings time (at the beginning of the running, the program prints 'DST' or 'Not DST').

Right now, I am focusing on making coherence plots in a spectrogram style (similar to the matlab 'coh_carpet' function) and a spectrogram depicting Gaussianity (similar to the plots made by the RayleighMonitor).  I have also been working on my  final paper and presentation.

  7191   Wed Aug 15 11:44:35 2012 jamieSummaryLSCntp installed on all workstations

Quote:

5) DTT wasn't working on rossa. Used the Date/Time GUI to reset the system time to match fb and then it stopped giving 'Test Timed Out'. Jamie check rossa ntpd.

ntp is now installed on all the workstations.  I also added it to the /users/controls/workstation-setup.sh script

  7190   Wed Aug 15 11:40:15 2012 YaakovSummarySTACISWeekly Summary

This week I've been focusing mainly on two things: 1) Designing a port for the STACIS that will allow external actuation and/or local feedback and 2) Investigating the seismic differential motion along the interferometer arms.

The circuit for the port is just a signal summing junction (in case we want to do feedforward and feedback at the same time) with BNC inputs for the external signal and switches that allow you to turn the external signal or feedback signal on/off. I'll test this on a breadboard and post the schematic if it works. I looked at the noise of the geophone pre-amp and DAC, which would be the feedback and external signal sources, respectively. According to Rolf Bork, the DAC noise is 700 nV/rtHz, and I measured the pre-amp board's minimum noise level at 20*10^-6 V/rtHz (which seems quite high). Both these noises are higher than the op-amp noise for my circuit (I'm considering the op-amp LT1012), which according to the specs is 30 nV/rtHz. This confirms that my circuit will not be the limiting noise source

Along with Den, I calibrated the seismometers in the lab and measured the displacement differential arm motion (see eLog 7186: http://nodus.ligo.caltech.edu:8080/40m/7186). I'm trying to find a transfer function for the seismic stacks (and pendulum, but that's simpler) so I can calculate the differential motion in the chamber. After doing this offline, I'll make new channels in the PEM to look at the ground and chamber differential motion along the arms online.

I also am looking at the noise of the geophones with their shunt resistor (4k resistor across the coil) removed, to see if it improves the noise at low frequencies. My motivation for this was that the geophone specs show a better V/m/s sensitivity at low frequencies when the shunt resistor is removed, so the actual signal may become larger than the internal noise at these frequencies.

  7189   Wed Aug 15 10:40:16 2012 DenUpdateCDSaa filters

The lack of AA filter for MCL signal is RFM model strongly disturbed entering to OAF signal

aa.png

  7188   Wed Aug 15 09:09:45 2012 jamieUpdateLSCLSC whitening triggers

Quote:

I'm ~30% of the way through implementing LSC whitening filter triggers.  I think that everything I have done should be compile-able, but please don't compile c1lsc tonight.  I haven't tested it, and some channel names have changed, so I need to fix the LSC screen when I'm not falling asleep.

Also, Rana pointed out that we may not want the whitening to trigger on immediately upon acquiring lock - if there are other modes ringing down in the cavity, or some weird transients, we don't want to amplify those signals.  We want to wait a second or so for them to die down, then turn on analog whitening.  Jamie - do you know how long the "unit delay" delays things in the RCG?  Do those do what I naively think they do?  I'll ask you in the morning.

The unit delay delays for a single cycle, so I think that's not what you want.  I'm not sure that there's an existing part to add delays like that.

We also need to be a little clever about it, though, since we'll want it to flip off if we loose lock during the delay.

  7187   Wed Aug 15 04:03:55 2012 ranaSummaryLSCY-Arm Locking

0) Did a bunch of alignment to get beams roughly centered on ETMY and ITMY and maximize power. Adjusted the aperture and focus on ETMY camera to get nice image. Camera needs to be screwed in tightly and cables given some real strain relief, Steve.

1) snapshots not working on many MEDM screens. Who's on top of this?

2) save/restore not working for PZT2 sliders

3) changed power and filter triggers on yarm to match xarm

4) yarm filters copied from xarm (need to handtune RGs)

5) DTT wasn't working on rossa. Used the Date/Time GUI to reset the system time to match fb and then it stopped giving 'Test Timed Out'. Jamie check rossa ntpd.

6) Removed the high 3.2 Hz RG filter. We don't have any sharp features like that in the spectrum.
   ---then added it back. The 3.2 Hz comes and goes depending on what Yoichi is doing over in the MC area. Leaving it in by default, but lowering the Q from 2 to 1.5.

7) Attached is the noise spectra, coherence, and loop gain model for this yarm condition. For the plant model, I assume a pendulum (f=1 Hz, Q = 9) and a cavity pole of 1600 Hz. Gain is scaled to set the UGF at 165 Hz (as guessed by looking at the servo gain peaking frequency). This cheezy model doesn't include any of the delays from DAC, AA, or AI. Eric and Sasha should have something more useful for us by Friday.

8) Change the DQ channels: need XARM and YARM IN1 at 16k. e.g. XARM_ERR, etc.

9) To get the DTT plots to make thumbnails in the elog, I print a .ps file and then use 'epstopdf' to make the PDF.

Attachment 1: yArmNoise_120815.png
yArmNoise_120815.png
Attachment 2: yarm.pdf
yarm.pdf
  7186   Wed Aug 15 01:14:19 2012 YaakovUpdatePEMDifferential Motion of X and Y Arm

Den and I measured the differential motion of the x and y arms using Guralp 1 at the end of the y arm, Guralp 2 at the beamsplitter, and the Streckeisen at the end of the x arm.

I calibrated the Streckeisen to the Guralp by calculating the relative gain of the seismometer signals at the microseism. The Guralp 1-y amplitude was 1.0237 times Guralp 2-y and Guralp 2-x was 38.54 times STS-x. The Guralp calibration (to go from counts to meters) I used was 0.61/1000/800/80/(2*pi*f) m/count.

The differential motion should keep decreasing at low frequencies because the ground will move together at such large wavelengths. It goes up because the seismometer noise begins to dominate at low frequencies (below about 0.5 Hz). Another possible error source could be that the seismometers are not perfectly aligned along the arm.

diff_motion_x_arm.pngdiff_motion_y_arm.png

Attachment 1: diff_motion_x_arm.png
diff_motion_x_arm.png
Attachment 2: diff_motion_y_arm.png
diff_motion_y_arm.png
  7185   Wed Aug 15 00:52:17 2012 DenUpdateWienerFilteringfilter calculation

A Matlab script to calculate Wiener filter coefficients and convert fir to iir is ready. Input is a file with zero mean witness and desired signals, output is a Foton zpk command to specify iir filter.

The plot shows comparison of offline fir , iir and online iir filtering. Spectrum below 4 Hz is still oscillating due to acoustic coupling, this is not a filtering effect. At 1 Hz actuator is badly compensated, more work should be done. Other then that online and offline filtering are the same.

wiener.png

  7184   Tue Aug 14 22:16:46 2012 JenneUpdateLSCLSC whitening triggers

I'm ~30% of the way through implementing LSC whitening filter triggers.  I think that everything I have done should be compile-able, but please don't compile c1lsc tonight.  I haven't tested it, and some channel names have changed, so I need to fix the LSC screen when I'm not falling asleep.

Also, Rana pointed out that we may not want the whitening to trigger on immediately upon acquiring lock - if there are other modes ringing down in the cavity, or some weird transients, we don't want to amplify those signals.  We want to wait a second or so for them to die down, then turn on analog whitening.  Jamie - do you know how long the "unit delay" delays things in the RCG?  Do those do what I naively think they do?  I'll ask you in the morning.

  7183   Tue Aug 14 21:01:51 2012 ranaUpdatePEMBLRMS

Screenshot-Untitled_Window.png

I fixed up the seismic.stp file for the StripTool display:

  1. All BLRMS channels now have a y-axis range of 3 decades. So they all are displaying the same relative changes.
  2. So the 0.01-0.1 Hz band which is all over the place is real, sort of. Masha says that it is due to the seismometer signal being dominated by noise below 0.1 Hz. She is going to fix this somehow.
  3. I changed the samping time from 1 sec. to 10 sec. to make the traces less fuzzy.
  4. We (Masha / Liz) should harmonize the colors of this file with what's on the summary pages.
  7182   Tue Aug 14 17:47:44 2012 JamieUpdateCDSc1sus machine replaced

Rolf and Alex came back over with a replacement machine for c1sus.   We removed the old machine, removed it's timing, dolphin, and PCIe extension cards and put them in the new machine.  We then installed the new machine and booted it and it came up fine.  The BIOS in this machine is slightly different, and it wasn't having the same failure-to-boot-with-no-COM issue that the previous one was.  The COM ports are turned off on this machine (as is the USB interface).

Unfortunately the problem we were experiencing with the old machine, that unloading certain models was causing others to twitch and that dolphin IPC writes were being dropped, is still there.  So the problem doesn't seem to have anything to do with hardware settings...

After some playing, Rolf and Alex determined that for some reason the c1rfm model is coming up in a strange state when started during boot.  It runs faster, but the IPC errors are there.  If instead all models are stopped, the c1rfm model is started first, and then the rest of the models are started, the c1rfm model runs ok.  They don't have an explanation for this, and I'm not sure how we can work around it other than knowing the problem is there and do manual restarts after boot.  I'll try to think of something more robust.

A better "fix" to the problems is to clean up all of our IPC routing, a bunch of which we're currently doing very inefficient right now.  We're routing things through c1rfm that don't need to be, which is introducing delays.  It particular, things that can communicate directly over RFM or dolphin should just do so.  We should also figure out if we can put the c1oaf and c1pem models on the same machine, so that they can communicate directly over shared memory (SHMEM).  That should cut down on overhead quite a bit.  I'll start to look at a plan to do that.

 

  7181   Tue Aug 14 16:33:51 2012 SashaUpdateComputer Scripts / ProgramsSimPlant indicator added

I added an indicator to the watch dog screen so that a little "SP" icon appears whenever the SimPlant is on. Since we only have one simplant (ETMX), only ETMX has the simPlant indicator. However, since assymetry is ugly, I moved all of the OL icons over so that they're in a line and so that there is room for future SP icons.

I also fixed the link to the Watchdogs on the main SUS screens (it was dead, but now it is ALIVE).

  7180   Tue Aug 14 16:19:12 2012 JenneUpdateGreen LockingXend doubling crystal heater unplugged, replugged

I went down to the Xend table to look at it to understand Steve's proposal, and I noticed that the doubling crystal's heater's cable is mushed between the table's edge and the black table cover wall.  This made me sad, so I disabled the heater, turned it off, then unplugged the cable from the back of the controller.  I tried to re-route the cable through the hole in the black table cover wall, but going that way the cable is ~1 foot too short.  So I put it back the way it was, but used a totally hacky solution to prevent the cable from being mushed.  I put a dog clamp right at the edge of the table so it is pushing on the table cover wall a little bit, to give the cable space to get out.  This is very mickey mouse, and kind of lame.  But we either need to make a cable extension, or somehow get the heater controller to sit much, much higher under the table.

I plugged the heater controller back in, and turned it back on to the same setpoint that it was at (I think 37.5C).  It's probably warm by now, but when I turned it back on, the heater's actual temp was 33C.

  7179   Tue Aug 14 15:58:44 2012 JenneUpdateGeneralTranslation to English: larger optical tables at the ends

Quote:

Quote:

I'm proposing larger optical tables at the ends to avoid the existing overcrowding. This would allow the initial pointing and optical level beams to set up correctly.

The existing table is 4 x 2 would be replaced by 4' x 3'   We would lose only ~3" space  toward exist door.

I'm working on the new ACRYLIC TABLE COVER for each end that will cost around $4k ea.  The new cover should fit the larger table.

Let me know what you think.

I'm not sure I see the motivation.  The tables are a little tight, but not that much.  If the issue is the incidence angle of the IP and OPLEV beams, then can't we solve that just by moving the table closer to the viewport?

The overcrowding alone doesn't seem bad enough to justify replacing the tables.

Steve pointed out to me (this is not in his original elog, although you can see it in the photo if you look closely), that we can't really move the table legs any closer to the chamber.  We have maybe 3" of clearance between the table leg and the blue support tube that supports the bottom of the stack.  Therefore, we can't just

So Steve's proposal is to leave the legs exactly where they are, and just put a larger table on those legs.  This leaves 9" unsupported on the chamber side, and 3" unsupported on the far side.  The tables are 4" thick. 

Steve also mentions that we will lose 1.5" on all 4 sides of the table, with the new acrylic boxes, so we'll be down to 1'9" unless we get the larger table, in which case we'd have 2'9", and 3'9" on the long direction.

I would like to see a sketch of the end tables, so we can see if 1'9" x 3'9" is enough.  Manasa is working on a new end table layout in parallel to the ringdown stuff.  If we're actually concerned about the input angle of the oplevs, then to fix that we need to either get the bigger table and hang it off the edge of the legs, or perhaps as Dmass suggested, get a "doggy cone collar", and give ourselves a larger opening angle of access to the viewport, from the current table location.

 

  7178   Tue Aug 14 14:26:40 2012 SteveUpdateCamerascameras touched up

 I optimized the TM views with illuminator light on quad1  It actually looks better there.

I'll post a dark-  OSEM light only in jpg tomorrow.  ETMY camera is malfunctioning in dark condition now.

 

Attachment 1: cameras.png
cameras.png
  7177   Tue Aug 14 13:21:34 2012 JenneUpdateCDSIPC senders no nothing about rates of IPC receivers, we need to filter signals appropriately

Quote:

Quote:

When signals are transmitted between the models running at different rates, no AI or AA filters are automatically applied. We need to fix our models.

This is known, but we just haven't fully groked it yet.  We need to look closely at every place we have IPCs between models running at different rates.  The sender has no information about receivers, so it can't reasonably do anything to pre-filter the signal on it's own.

So for transmission from:

  • faster -> slower models: add anti-aliasing on the sender side
  • slower -> faster models: add anti-imaging on the receiver side

 *sigh* This is one of those things that I meant to take care of months ago, but haven't yet.  I agree that it needs doing.  It's been on my whiteboard to-do list for a long time now.  Bad Jenne for not taking care of it.

  7176   Tue Aug 14 11:49:15 2012 DenUpdateCDSDebugging of c1sus machine and c1rfm models

Quote:

 

  We might just need to reduce the load on c1rfm, maybe by introducing a c1rfm2?

 

 A huge data flow goes from PEM to OAF through RFM. I think we need to make PEM and OAF run on the same machine and transmit signals through the shared memory.

ELOG V3.1.3-